# ImpMIA: Leveraging Implicit Bias for Membership Inference Attack under Realistic Scenarios

### v1

@@ Abstract
Determining which data samples were used to train a model—known as Membership Inference Attack (MIA)—is a well-studied and important problem with implications for data privacy. Black-box methods presume access only to the model’s outputs and often rely on training auxiliary reference models. While they have shown strong empirical performance, they rely on assumptions that rarely hold in real-world settings: (i) the attacker knows the training hyperparameters; (ii) all available non-training samples come from the same distribution as the training data; and (iii) the fraction of training data in the evaluation set is known. In this paper, we demonstrate that removing these assumptions leads to a significant drop in the performance of black-box attacks. We introduce ImpMIA, a Membership Inference Attack that exploits the Implicit Bias of neural networks, hence removes the need to rely on any reference models and their assumptions. ImpMIA is a white-box attack – a setting which assumes access to model weights and is becoming increasingly realistic given that many models are publicly available (e.g., via Hugging Face). Building on maximum-margin implicit bias theory, ImpMIA uses the Karush–Kuhn–Tucker (KKT) optimality conditions to identify training samples. This is done by finding the samples whose gradients most strongly reconstruct the trained model’s parameters. As a result, ImpMIA achieves state-of-the-art performance compared to both black and white box attacks in realistic settings where only the model weights and a superset of the training data are available.
@@ 1 Introduction
Ensuring that trained models do not leak information about their training sets is a critical challenge. Membership inference attacks (MIAs) evaluate this risk by determining whether a given example was part of a model’s training data. MIAs can be broadly divided into two categories: black-box, which assume only query access to model outputs (Shokri et al., 2017; Yeom et al., 2018; Li & Zhang, 2021; Carlini et al., 2022), and white-box, which exploit access to internal parameters such as weights or gradients (Nasr et al., 2019; Leino & Fredrikson, 2020; Cohen & Giryes, 2024).
The most effective black-box MIAs are reference-model–based attacks. These methods estimate the distribution of losses for members (training samples) versus non-members by training auxiliary reference models that mimic the target model, thereby learning its loss behavior. However, training large sets of reference models is computationally expensive, and—more importantly—their effectiveness depends on the reference models being accurate surrogates of the target. As a result, these attacks rely on several strong assumptions: (i) the attacker knows the training hyperparameters (e.g., learning rate, optimizer, number of epochs); (ii) the non-training samples come from the same distribution as the training data; and (iii) the fraction of training members in the evaluation set is known. When any of these assumptions is violated, the performance of black-box MIAs drops significantly (Shokri et al., 2017; Salem et al., 2019; Song & Mittal, 2021; Carlini et al., 2022), limiting their reliability for auditing privacy in realistic settings.
On the other hand, white-box MIAs assume access to model weights or gradients. While this is still an assumption, this scenario is increasingly realistic, as many modern models are released with their full parameters publicly available (e.g., via platforms such as Hugging Face). White-box MIAs have shown strong performance by leveraging gradients and activations (Nasr et al., 2019; Leino & Fredrikson, 2020), or influence scores (Cohen & Giryes, 2024). However, despite these advances, current white-box attacks still fall short compared to the best black-box approaches when evaluated under stringent criteria such as the true-positive rate (TPR) at very low false-positive rates (FPR), which has been suggested as a reliable evaluation perspective by Carlini et al. (2022).
We propose ImpMIA, a white-box membership inference attack that is the first to adapt neural network implicit bias theory for this task (see our approach overview in Figure 1). Unlike prior approaches, our attack does not rely on reference models and is therefore unaffected by the common assumptions in reference-model methods. ImpMIA requires no knowledge of the target model’s training procedure or data distribution, operating in a realistic scenario where only the model weights and a superset of the training data are available.
@@
Our attack builds on the theory of implicit bias in neural networks, which shows that gradient-based optimization tends to converge to solutions that satisfy the Karush–Kuhn–Tucker (KKT) optimality conditions of a certain maximum-margin problem (Lyu & Li, 2019; Ji & Telgarsky, 2020). In practice, this implies that the trained parameters of a network can be approximately expressed as a linear combination of per-sample gradients from the training set. Given a set of candidate samples and the trained network weights, we optimize a set of coefficients—one for each sample—that best reconstructs the network parameters. This provides the key signal: training samples are expected to receive significantly larger coefficients, while non-members remain small.
ImpMIA achieves superior results, surpassing both black-box and white-box attacks under realistic settings across three benchmark datasets. Our extensive analysis shows that removing the knowledge assumptions made by most reference model based methods, leads to a significant drop in the performance of state-of-the-art methods, while our attack remains unaffected. Altogether, these results highlight the importance and effectiveness of our proposed method as a practical membership inference attack in realistic scenarios, likely to be most relevant for real-world applications.
@@
  • ImpMIA achieves state-of-the-art performance in realistic scenarios where only model weights and a candidate data pool are available.
  • We provide a systematic evaluation of the robustness of state-of-the-art MIA methods under realistic conditions where training hyperparameters, data distribution, or member ratios are unknown.
@@ 2.1 Membership inference
Membership Inference Attacks (MIAs) are typically divided into black-box attacks, which rely only on model outputs, and white-box attacks, which exploit access to model parameters.
@@
White-box attacks exploit access to a model’s internal parameters (often gradients) to amplify membership signals. Nasr et al. (2019) introduced one of the first frameworks, leveraging activations and per-example gradients. Sablayrolles et al. (2019) derived the Bayes-optimal test under white-box access, showing that maximum membership power can be achieved by computing likelihood ratios over model parameters. Leino & Fredrikson (2020) showed in their Stolen Memories attack that gradient norms alone provide strong membership signals, and that training an auxiliary classifier to distinguish gradients from members versus non-members further improves accuracy. Most recently, Cohen & Giryes (2024) proposed a self-influence attack that uses influence functions to measure each sample’s effect on its own loss, combined with the predicted label. While white-box access is a strong assumption, it is increasingly realistic as many modern models are released with their full weights (e.g., on Hugging Face).
@@
Black-box MIAs assume the attacker can only query the target model and observe its outputs. Shokri et al. (2017) introduced the shadow-model framework, training reference models to mimic the target and then learning an attack model from their outputs. Yeom et al. (2018) later showed that even without reference models, the simple “gap” heuristic—predicting membership when the model’s output label matches the ground truth, can be effective. Li & Zhang (2021) proposed decision-based attacks relying on adversarial perturbations, while Ye et al. (2022) introduced Attack-P, a population-based loss thresholding method, and Attack-R, a sample-specific calibration using percentiles from reference models for improved robustness. Building on the reference-model paradigm, Carlini et al. (2022) proposed LiRA, which compares target losses to reference-model distributions and emphasized low false-positive evaluation. Zarifzadeh et al. (2023) (RMIA) further refined LiRA with an optimized likelihood-ratio test, improving efficiency under strict computational limits. LiRA and RMIA currently represent the strongest-performing black-box MIAs.
@@
Limitations of Reference-Model Attacks. Reference-model attacks are costly to train and their effectiveness depends heavily on how these models are trained. Specifically, three key assumptions which violating them significantly reduces MIA performance: (i) Knowledge of the target model’s training hyperparameters (e.g., learning rate, optimizer, epochs): Jayaraman et al. (2020) and Carlini et al. (2022) (LiRA) reported accuracy drops when altering those hyperparameters. (ii) Matching data distribution: Salem et al. (2019) shows degraded accuracy when shadows were trained on different domains (e.g., CIFAR-10 for a CIFAR-100 target). (iii) Member ratio in the inference pool: both Jayaraman et al. (2020) and Song & Mittal (2021) found inflated false positives and reduced accuracy when this assumption was wrong. In this work, we systematically test these three factors showing significant drops across models on CIFAR-10, CIFAR-100, and CINIC-10. Such scenarios—where these factors are unknown to the attacker—are more realistic, as most models, particularly those trained on sensitive data, are unlikely to publish this information.
@@ 2.2 Implicit Bias of Gradient Descent
In overparameterized neural networks, one might expect overfitting the training data. Yet gradient-based methods tend to converge to classifiers that generalize well to new unseen data (Zhang et al., 2021; Neyshabur et al., 2017). This phenomenon is explained by the implicit bias of training algorithms: gradient descent tends to prefer specific solutions, and characterizing these has been central to deep-learning theory in recent years (see Vardi (2023) for a survey). For homogeneous ReLU networks trained to zero logistic or cross-entropy loss, Lyu & Li (2019) and Ji & Telgarsky (2020) showed that the learned weights necessarily satisfy the KKT conditions of a maximum-margin problem. Building on this, Haim et al. (2022) demonstrated that networks trained with binary cross-entropy allow reconstruction of dozens of nearly pixel-perfect training samples. This was later extended to multiclass classifiers (Buzaglo et al., 2023) and more realistic transfer-learning workflows (Oz et al., 2024). All of these data reconstruction attacks are limited to very small datasets (up to a few thousand examples), and require simple models (mostly MLPs). In this work, we adapt the implicit-bias approach for the first time to membership inference attack, identifying which samples from a candidate pool best satisfy the KKT conditions.
@@ 3 Preliminaries
3.1 Membership Inference Attacks Setting
Membership inference attacks (MIAs) aim to determine which data points were part of the training set of a machine learning model. The strongest recent black-box methods are reference-model based, and they rely on additional assumptions about the target model’s training configuration and the candidate set, which are unlikely to hold in practice. In this work, we focus on a realistic scenario where the adversary has access only to a superset that contains the training set and to the model weights. This reflects real-world conditions: modern models are often released publicly with their weights, while auditors may possess large candidate pools that include the training set but lack detailed knowledge of the training data distribution or the exact training configuration. Our setting adapts the basic membership inference game (Yeom et al., 2018; Jayaraman et al., 2020), where a single sample is evaluated at a time, into a superset-based formulation in which the full candidate pool is attacked and evaluated (similar to the online setting in Carlini et al. (2022); Zarifzadeh et al. (2023)). We assume that the superset contains the full training set or at least a large portion of it. While this assumption was not taken in prior work, the reported results in those settings were in fact obtained under it. That is, for technical reasons, their evaluations were obtained in a setting where the attacker uses a superset of the training data (Carlini et al., 2022). Results for our method in a scenario where this assumption does not hold are presented in Appendix B.3.
Formally, we study an assumption-free superset setting. Let $X_{\mathrm{train}}$ be the (unknown) training set drawn from a distribution $\mathcal{D}$, and $f_{\theta}$ a model trained on $X_{\mathrm{train}}$. Let $X_{\mathrm{sup}}$ be a candidate pool such that $X_{\mathrm{train}}\subseteq X_{\mathrm{sup}}$, with the remaining samples being non-members, potentially drawn from other distributions. The adversary is given the trained parameters $\theta$ and the pool $X_{\mathrm{sup}}$, but: (i) does not know the hyperparameters used to train the target model; (ii) cannot assume non-members are drawn from the same distribution as the training set; and (iii) does not know how many members are in $X_{\mathrm{sup}}$ or their ratio. The adversary must then assign a real-valued score to each sample, with membership decisions.
@@ 3.2 The Implicit Bias Formulation
In this section, we provide an overview of the KKT conditions and the maximum-margin formulation, following the definitions in Haim et al. (2022); Buzaglo et al. (2023). While the theory described below is formally constrained to homogeneous [^1] A model $\Phi$ is homogeneous w.r.t. $\theta$ if there exists $L>0$ such that $\forall c>0,x:\;\Phi(x;c\theta)=c^{L}\Phi(x;\theta)$. ReLU networks (Lyu & Li, 2019; Ji & Telgarsky, 2020), we show in practice that the results hold more generally for other architectures. The theoretical results of Lyu & Li (2019); Ji & Telgarsky (2020) consider training without weight decay, but in Appendix D we show that incorporating weight decay leads to the same final equations, and in Section B.5 we analyze its influence on the attack performance.
Implicit bias of gradient flow on homogeneous networks: Let $\Phi(\theta;\cdot):\mathbb{R}^{d}\to\mathbb{R}^{C}$ be a homogeneous ReLU network with outputs $\Phi_{j}(\theta;x)$. Consider minimizing the cross-entropy loss over a multiclass dataset $\{(x_{i},y_{i})\}_{i=1}^{n}\subseteq\mathbb{R}^{d}\times[C]$ using gradient flow (i.e., gradient descent with a small step size). Suppose that at some time $t_{0}$ the network classifies all training samples correctly. Then gradient flow converges in direction to a KKT point of the multiclass maximum-margin problem:
$\min_{\theta}\;\tfrac{1}{2}\|\theta\|^{2}\quad\text{s.t.}\quad\Phi_{y_{i}}(\theta;x_{i})-\Phi_{j}(\theta;x_{i})\geq 1,\quad\forall i\in[n],\ \forall j\in[C]\setminus\{y_{i}\}.$
The corresponding KKT conditions for solving this maximum margin problem are:
$\theta-\sum_{i\in[n]}\sum_{j\in[C]\setminus\{y_{i}\}}\lambda_{i,j}\,\nabla_{\theta}\bigl[\Phi_{y_{i}}(\theta;x_{i})-\Phi_{j}(\theta;x_{i})\bigr]=0,$ (1)
$\forall\,i\in[n],\;j\in[C]\setminus\{y_{i}\}:\;\begin{cases}\Phi_{y_{i}}(\theta;x_{i})-\Phi_{j}(\theta;x_{i})\;\geq\;1,\\[1.0pt] \lambda_{i,j}\;\geq\;0,\\[1.0pt] \lambda_{i,j}=0\;\;\text{if }\Phi_{y_{i}}(\theta;x_{i})-\Phi_{j}(\theta;x_{i})\neq 1.\end{cases}$ (2)
Equation 1, called the stationarity condition, represents the weights as a linear combination of margin gradients (distance between a sample’s true class $\Phi_{y_{i}}$ and other classes $\Phi_{j}$), while Equation 2 specifies additional constraints. The coefficients are the $\lambda_{i,j}$, defined per sample i per class j. In practice, the distance of a sample $x_{i}$ to the decision boundary is typically determined by a single competing class $j$. Therefore, following Buzaglo et al. (2023), we simplify the first condition by
considering only the class with the smallest margin:
$\theta\;=\;\sum_{i=1}^{n}\lambda_{i}\,g_{i},\ \ \ \ g_{i}=\nabla_{\theta}\Big[\Phi_{y_{i}}(x_{i};\theta)-\max_{j\neq y_{i}}\Phi_{j}(x_{i};\theta)\Big].$ (3)

### v3

@@ Abstract
Determining which data samples were used to train a model, known as Membership Inference Attack (MIA), is a well-studied and important problem with implications on data privacy. SotA methods (which are black-box attacks) rely on training many auxiliary reference models to imitate the behavior of the attacked model. As such, they rely on assumptions which rarely hold in real-world settings: (i) the attacker knows the training hyperparameters; (ii) all available non-training samples come from the same distribution as the training data; and (iii) the fraction of training data in the evaluation set is known. We show that removing these assumptions significantly harms the performance of black-box attacks. We introduce ImpMIA, a Membership Inference Attack that exploits the Implicit Bias of neural networks. Building on the maximum-margin implicit bias theory, ImpMIA uses the Karush–Kuhn–Tucker (KKT) optimality conditions to identify training samples – those whose gradients most strongly reconstruct the trained model’s parameters. Our approach is optimization-based, and requires NO training of reference-models, thus removing the need for any knowledge/assumptions regarding the attacked model’s training procedure. While ImpMIA is a white-box attack (a setting which assumes access to model weights), this is becoming increasingly realistic given that many models are publicly available (e.g., via Hugging Face). ImpMIA achieves SotA performance compared to both black and white box attacks in settings where only the model weights are known, and a superset of the training data is available.
@@ 1 Introduction
Ensuring that trained models do not leak information about their training sets is a critical challenge. Membership inference attacks (MIAs) evaluate this risk by determining whether a given example was part of a model’s training data. MIAs can be broadly divided into two categories: black-box, which assume only query access to model outputs (Shokri et al., 2017; Yeom et al., 2018; Li and Zhang, 2021; Carlini et al., 2022; Wu et al., 2024; Peng et al., 2024), and white-box, which exploit access to internal parameters such as weights or gradients (Nasr et al., 2019; Leino and Fredrikson, 2020; Cohen and Giryes, 2024; Suri et al., 2024).
The most effective black-box MIAs are reference-model-based attacks. These methods estimate the distribution of losses for members (training samples) versus non-members by training auxiliary reference models that mimic the target model, thereby learning its loss behavior. However, training large sets of reference models is computationally expensive, and—more importantly—their effectiveness depends on the reference models being accurate surrogates of the target. As a result, these attacks rely on several strong assumptions: (i) the attacker knows the training hyperparameters (e.g., learning rate, optimizer, number of epochs); (ii) the non-training samples come from the same distribution as the training data; and (iii) the fraction of training members in the evaluation set is known. When any of these assumptions is violated, the performance of black-box MIAs drops significantly (Shokri et al., 2017; Salem et al., 2019; Song and Mittal, 2021; Carlini et al., 2022), limiting their reliability for auditing privacy.
On the other hand, white-box MIAs assume access to model weights or gradients. While this is a restrictive assumption, this scenario is increasingly realistic, as many modern models are released with their parameters publicly available (e.g., via platforms such as Hugging Face). White-box MIAs have shown strong performance by leveraging gradients and activations (Nasr et al., 2019; Leino and Fredrikson, 2020), or influence scores (Cohen and Giryes, 2024). However, despite these advances, current white-box attacks still lag behind the best black-box approaches when evaluated under stringent criteria such as the true-positive rate (TPR) at very low false-positive rates (FPR), which has been suggested as a reliable evaluation criterion by Carlini et al. (2022).
We propose ImpMIA, a white-box membership inference attack that is the first to adapt neural network Implicit-Bias theory for this task (see Approach Overview in Fig. 1). Unlike prior approaches, our attack does not rely on training any reference models, and is therefore unaffected by the common assumptions in reference-model methods. ImpMIA requires no knowledge of the target model’s training procedure or data distribution, operating in a no auxiliary-knowledge scenario where only the model weights are known, and a superset of the training data is available.
@@
Our attack builds on the implicit-bias theory in neural nets, which shows that gradient-based optimization tends to converge to solutions that satisfy the Karush–Kuhn–Tucker (KKT) optimality conditions of a certain maximum-margin problem (Lyu and Li, 2019; Ji and Telgarsky, 2020). In practice, this implies that the trained parameters of a network can be approximately expressed as a linear combination of per-sample gradients from the training set. Given a set of candidate samples and the trained network weights, we optimize a set of coefficients—one per sample—that best reconstructs the network parameters. This provides the key signal: training samples are expected to receive significantly larger coefficients, while non-members remain small.
ImpMIA is best suited for scenarios where the training-set is contained in the examined-set. Such scenarios are relevant, e.g., if one wants to check whether their own dataset has been used for training a model, or when a dataset is known to contain the training-set (e.g., a publicly available dataset), but not which part of it was used for training. While this is a somewhat restrictive assumption, we show that it can be relaxed (App. B.3), and that our performance is competitive even at low coverage of the training-set (only 10%), and becomes superior at high coverage.
ImpMIA achieves superior results, surpassing both black-box and white-box attacks under no auxiliary-knowledge settings across 3 benchmark datasets. Our extensive analysis shows that removing the knowledge assumptions made by most reference-model-based methods (knowledge of the attacked model’s training procedure & setting), leads to a significant drop in performance of SotA methods, while our attack remains unaffected. Altogether, these results highlight the importance and effectiveness of our proposed method as a practical membership inference attack.
@@
  • ImpMIA achieves SotA performance in scenarios where only model weights & a candidate data pool are available.
  • We provide a systematic evaluation of the robustness of leading MIA methods in settings where training hyperparameters, data distribution, or member ratios are unknown.
@@ 2.1 Membership inference
Membership Inference Attacks (MIAs) are divided into black-box attacks, which rely on model outputs, and white-box attacks, which exploit access to model parameters.
@@
White-box attacks exploit access to a model’s internal parameters (often gradients) to amplify membership signals. Nasr et al. (2019) introduced one of the first frameworks, leveraging activations and per-example gradients. Sablayrolles et al. (2019) derived the Bayes-optimal test under white-box access, showing that maximum membership power can be achieved by computing likelihood ratios over model parameters. Leino and Fredrikson (2020) showed in their Stolen Memories attack that gradient norms alone provide strong membership signals, and that training an auxiliary classifier to distinguish gradients from members versus non-members further improves accuracy. Most recently, Cohen and Giryes (2024) proposed a self-influence attack that uses influence functions to measure each sample’s effect on its own loss, combined with the predicted label. While white-box access is a strong assumption, it is increasingly realistic as many modern models are released with their full weights (e.g., on Hugging Face).
@@
Black-box MIAs assume the attacker can only query the target model and observe its outputs. Shokri et al. (2017) introduced the shadow-model framework, training reference models to mimic the target and then learning an attack model from their outputs. Yeom et al. (2018) later showed that even without reference models, the simple “gap” heuristic, predicting membership when the model’s output label matches the ground truth, can be effective. Li and Zhang (2021) proposed decision-based attacks relying on adversarial perturbations, while Ye et al. (2022) introduced Attack-P, a population-based loss thresholding method, and Attack-R, a sample-specific calibration using percentiles from reference models for improved robustness. Building on the reference-model paradigm, Carlini et al. (2022) proposed LiRA, which compares target losses to reference-model distributions and emphasized low false-positive evaluation. Zarifzadeh et al. (2023) (RMIA) further refined LiRA with an optimized likelihood-ratio test, improving efficiency under strict computational limits. LiRA and RMIA currently represent the strongest-performing black-box MIAs.
@@
Limitations of Reference-Model Attacks. Reference-model attacks are costly to train and their effectiveness depends heavily on how these models are trained. Specifically, three key assumptions; whose violation significantly reduces MIA performance: (i) Knowledge of the target model’s training hyperparameters (e.g., learning rate, optimizer, epochs): Jayaraman et al. (2021) and Carlini et al. (2022) (LiRA) reported accuracy drops when altering those hyperparameters, while Pradhan et al. (2025) propose methods to partially mitigate this hyperparameters sensitivity, but rely on additional reference models and distributional assumptions. (ii) Matching data distribution: Salem et al. (2019) and Shi et al. (2024) show degraded accuracy when shadows were trained on different domains (e.g., CIFAR-10 for a CIFAR-100 target). (iii) Member ratio in the inference pool: Jayaraman et al. (2021) and Song and Mittal (2021) found inflated false positives and reduced accuracy when this assumption was wrong. In this work, we systematically test these factors showing significant drops across models on CIFAR-10, CIFAR-100, and CINIC-10. Such scenarios where these factors are unknown to the attacker are common as models, particularly those trained on sensitive data, are unlikely to publish this information.
@@ 2.2 Implicit Bias of Gradient Descent
In overparameterized neural networks, one might expect overfitting the training data. Yet gradient-based methods tend to converge to classifiers that generalize well to new unseen data (Zhang et al., 2021; Neyshabur et al., 2017). This phenomenon is explained by the implicit bias of training algorithms: gradient descent tends to prefer specific solutions, and characterizing these has been central to deep-learning theory in recent years (see Vardi (2023) for a survey). For homogeneous ReLU networks trained to zero logistic or cross-entropy loss, Lyu and Li (2019) and Ji and Telgarsky (2020) showed that the learned weights necessarily satisfy the KKT conditions of a maximum-margin problem. Building on this, Haim et al. (2022) demonstrated that networks trained with binary cross-entropy allow reconstruction of dozens of nearly pixel-perfect training samples. This was later extended to multiclass classifiers (Buzaglo et al., 2023) and more realistic transfer-learning workflows (Oz et al., 2024). All of these data reconstruction attacks are limited to very small datasets (up to a few thousand examples), and require simple models (mostly MLPs). In this work, we adapt the implicit-bias approach for the first time to membership inference attack, identifying which samples from a candidate pool best satisfy the KKT conditions.
@@ 3 Preliminaries
3.1 Membership Inference Attack Setting
Membership inference attacks (MIAs) aim to determine which data points were part of the training set of a machine learning model. The strongest recent black-box methods are reference-model based, and they rely on additional assumptions about the target model’s training configuration and the candidate set, which are unlikely to hold in practice. In this work, we focus on a scenario where the adversary has access only to a superset that contains the training set (or part of the training set) and to the model weights. This reflects real-world conditions: modern models are often released publicly with their weights, while auditors may possess large candidate pools that include the training set but lack detailed knowledge of the training data distribution or the exact training configuration. Our setting adapts the basic membership inference game (Yeom et al., 2018; Jayaraman et al., 2021), where a single sample is evaluated at a time, into a set-based formulation in which the full candidate pool is attacked and evaluated (similar to the online setting in Carlini et al. (2022); Zarifzadeh et al. (2023)). We assume that the evaluated set contains the training set or at least a portion of it (e.g. more than 10%). While this assumption was not taken in prior work, the reported results in those settings were in fact obtained under it. That is, for technical reasons, their evaluations were obtained in a setting where the attacker uses a superset of the training data (Carlini et al., 2022). Results for our method in a scenario where this assumption is relaxed are presented in Sec. B.3.
Formally, we study a set-based No-Auxiliary-Knowledge setting. Let $X_{\mathrm{train}}$ be the (unknown) training set drawn from a distribution $\mathcal{D}$, and $f_{\theta}$ a model trained on $X_{\mathrm{train}}$. Let $X_{\mathrm{sup}}$ be a candidate pool with $X_{\mathrm{train}}\subseteq X_{\mathrm{sup}}$, and the remaining samples being non-members, potentially drawn from other distributions. The adversary is given the trained parameters $\theta$ and the pool $X_{\mathrm{sup}}$, but: (i) does not know the target model’s training hyperparameters; (ii) cannot assume non-members are drawn from the same distribution as the training set; and (iii) does not know how many members are in $X_{\mathrm{sup}}$ or their ratio. The adversary must then assign a real-valued score to each sample, with membership decisions.
@@ 3.2 The Implicit Bias Formulation
In this section, we provide an overview of the KKT conditions and the maximum-margin formulation, following the definitions in Haim et al. (2022); Buzaglo et al. (2023). While the theory described below is formally constrained to homogeneous [^1] A model $\Phi$ is homogeneous w.r.t. $\theta$ if there exists $L>0$ such that $\forall c>0,x:\;\Phi(x;c\theta)=c^{L}\Phi(x;\theta)$. ReLU networks (Lyu and Li, 2019; Ji and Telgarsky, 2020), we show in practice that the results hold more generally for other architectures. The theoretical results of Lyu and Li (2019); Ji and Telgarsky (2020) consider training without weight decay, but in Appendix D we show that incorporating weight decay leads to the same final equations, and in Sec. B.6 we analyze its influence on the attack performance.
Implicit bias of gradient flow on homogeneous networks: Let $\Phi(\theta;\cdot):\mathbb{R}^{d}\to\mathbb{R}^{C}$ be a homogeneous ReLU network with logits $\Phi(\theta;x)$. Consider minimizing the cross-entropy loss over a multiclass dataset $\{(x_{i},y_{i})\}_{i=1}^{n}\subseteq\mathbb{R}^{d}\times[C]$ using gradient flow (i.e., gradient descent with a small step size). Suppose that at some time $t_{0}$ the network classifies all training samples correctly. Then gradient flow converges in direction to a KKT point of the multiclass maximum-margin problem:
$\min_{\theta}$ $\tfrac{1}{2}\|\theta\|^{2}$ (1)
s.t. $\Phi_{y_{i}}(\theta;x_{i})-\Phi_{j}(\theta;x_{i})\geq 1,$
$\forall i\in[n],\ \forall j\in[C]\setminus\{y_{i}\}.$
The KKT conditions for the multiclass maximum-margin problem are defined in terms of the pairwise margins. For each sample $i$ and competing class $j\neq y_{i}$, define
$m_{i,j}(\theta):=\Phi_{y_{i}}(\theta;x_{i})-\Phi_{j}(\theta;x_{i}).$ (2)
The KKT conditions are then given by:
$\theta=\sum_{i\in[n]}\sum_{j\in[C]\setminus\{y_{i}\}}\lambda_{i,j}\,\nabla_{\theta}m_{i,j}(\theta).$ (3)
$\text{For all }i\text{ and }j\neq y_{i}:$ $m_{i,j}(\theta)\geq 1,$ (4)
$\lambda_{i,j}\geq 0,$
$\lambda_{i,j}=0\text{ if }m_{i,j}(\theta)\neq 1.$
Equation 3, called the stationarity condition, represents the weights as a linear combination of margin gradients (distance between a sample’s true class $\Phi_{y_{i}}$ and other classes $\Phi_{j}$), while Equation 4 specifies additional constraints. The coefficients are the $\lambda_{i,j}$, defined per sample i per class j. In practice, the distance of a sample $x_{i}$ to the decision boundary is typically determined by a single competing class $j$. Following Buzaglo et al. (2023), we therefore simplify the stationarity condition by retaining only the smallest margin:
$m_{i}^{\min}(\theta)$ $=\Phi_{y_{i}}(x_{i};\theta)-\max_{j\neq y_{i}}\Phi_{j}(x_{i};\theta),$ (5)
$\theta$ $=\sum_{i=1}^{m}\lambda_{i}\,g_{i},\ \ \ \ g_{i}=\nabla_{\theta}m_{i}^{\min}(\theta).$